Can I get help with this problem please.
f(x)=5x+4; g(x)=7x-1 and solve for (f/g)(1)
I started with this 5x+4/7x-1 and got 5/7+33/7(7x-1) and that was wrong; just confused I believe on what they wanted.

Answer by Alan3354(69443)

f(x)=5x+4; g(x)=7x-1 and solve for (f/g)(1)
---
I started with this 5x+4/7x-1 and got 5/7+33/7(7x-1) and that was wrong;
I'm not sure about that division, might be right.
---
It's simpler to sub the 1 first and find f(1)/g(1).
= (5*1 + 4)/(7*1 - 1)
= 9/6
= 3/2